The position of an object moving in a straight line is given by s=t^2-8t, where s is in feet and t is the time in seconds the object has been in motion. How long (to the nearest tenth) will it take the object to move 10 feet?
Answer by stanbon(75887)

s=t^2-8t
10=t^2-8t
t^2-8t-10=0
Use the quadratic formula to get:
t=[8+-sqrt(64+40)]/2
t=[8+-2sqrt(26)]]2
t=[4+-sqrt26]
t= 9.09 seconds
Cheers,
Stan H.